# Comparative Analysis of Weather-Based Indexes and the Actuaries Climate IndexTM for Crop Yield Prediction

### v1

@@ Abstract
Climate change poses significant challenges to agriculture sector, affecting both crop productivity and economic stability. This study investigates the utility of climate indexes in predicting crop yields under diverse climatic conditions. Using generalized statistical models and machine learning algorithms, this research compares the predictive performance of long-established weather-based indexes with the Actuaries Climate IndexTM (ACI), a newer entrant in the field, showcasing its potential for real-time applications and broader applicability across various sectors. To enhance model reliability and address multicollinearity among weather-related variables, the study incorporates both principal component analysis and functional principal component analysis. A total of 22 models, each constructed with different explanatory variables, highlight the significant impact of wind speed and sea-level changes alongside temperature and precipitation on crop yield variability. Furthermore, the ACI exhibits strong predictive capabilities, offering valuable insights for climate risk assessment and decision-making in agriculture and insurance-related sectors.
Keywords actuaries climate index  $\cdot$ principal component analysis  $\cdot$ crop yield  $\cdot$ weather-based indexes  $\cdot$ machine learning
@@ 1 Introduction
Climate change refers to long-term shifts in temperature and weather patterns, primarily driven by human-induced greenhouse gas emissions. Recent reports from the Intergovernmental Panel on Climate Change highlight the significant impact of human activities on global temperatures over the past few decades (IPCC 2022; Masson-Delmotte et al. 2021). One of the most concerning consequences of climate change is the increase in the frequency and intensity of heatwaves, which exacerbate heat-related illnesses and deaths. WHO (2014) projects that climate change could result in an additional 250,000 deaths annually between 2030 and 2050. Historical events underline the severity of this issue. For instance, in August 2003, France endured its hottest summer in 50 years, leading to over 14,000 fatalities (Kovats et al. 2004). Similarly, a 2010 heatwave in Quebec, Canada, caused a 33% spike in daily mortality rates and hospital emergency visits (Bustinza et al. 2013).
The financial risks for the insurance industry is heightened by climate change, primarily due to the increasing frequency and severity of extreme weather events such as hurricanes, floods, and wildfires. Gall et al. (2011) highlight that direct losses from natural hazards in the United States (U.S.) are steadily increasing and could reach \$300 to \$400 billion within a single decade. A specific example is the California wildfires of 2018, which Wang et al. (2021) estimated to have caused \$148.5 billion in damages, including capital losses, health costs, and indirect economic disruptions. More recently, the wildfires in the Los Angeles area have the potential to inflict up to \$250 billion in economic losses, making them one of the most devastating natural disasters in U.S. history (Danielle 2025). From a broader perspective, Zhou et al. (2024) discuss the profound impact of climate change across various insurance sectors, including property/casualty, health, life, and crop insurance. They underscore the urgency for standardized methodologies to measure and quantify climate risks, ensuring that the industry can adapt to the sustainability challenges posed by a rapidly changing climate.
Climate change also leads to challenges like reduced crop yields and water scarcity, posing significant threats to global food security. Research by Lobell and Gourdji (2012) predicts that global warming could lower global crop yields by approximately 1.5% per decade, with potential reductions reaching as high as 4%. Several other studies reach similar results. For instance, Hu et al. (2024) observed that a 1°C rise in global temperatures results in a 6-8% average decrease in yields of major crops worldwide. Cai et al. (2009) project a 23-24% reduction in corn yields in central Illinois, U.S., by 2055, largely driven by increased water stress. Similarly, Adhikari et al. (2015) examine the impact of climate change on crops in Eastern Africa, concluding that the yields of major grains are significantly affected negatively, with wheat being particularly vulnerable—up to 72% of current yields are projected to decline.
Climate impact is mostly measured and interpreted by indexes. In the literature, various indexes such as the Climate Extremes Index (Karl et al. 1996), Greenhouse Climate Response Index (Karl et al. 1996), Palmer Drought Severity Index (Palmer 1965), Cooling Degree Days (Jewson and Brix 2005), Heating Degree Days (Jewson and Brix 2005) and Cumulative Rainfall Index (Odening et al. 2007) have been developed to quantify climate change, enabling researchers and policymakers to assess systematically its impact. These indexes often focus on critical variables such as temperature extremes, amounts of precipitation, drought severity, summarizing them into measurable indicators that facilitate the analysis of climate patterns and their consequences.
@@
In recent years, the Actuaries Climate IndexTM (ACI) has emerged as a pivotal tool for understanding and monitoring climate risks, offering valuable insights for insurance companies, governments, and the general public. Developed in 2016 by the four major North American actuarial societies, the ACI (2018) uses historical meteorological data from the United States and Canada to track extreme conditions in air temperature, precipitation, drought, wind, and sea-level rise. Its practical methodology has quickly been adapted for use in other countries. Curry (2015) demonstrates that the ACI methodology is applicable to regions like the United Kingdom and Europe, provided appropriate and sufficient meteorological data are available. The Australian Actuaries Climate Index (AACI 2018), which is tailored to Australian meteorological conditions, was introduced in 2018. The Turkish Actuaries Climate Index was developed in 2022 by Nevruz et al. (2022), and focuses on climate patterns specific to the Ankara region in Türkiye. Similarly, Zhou et al. (2023) launched the Iberian Actuarial Climate Index in 2023, marking the first application of the ACI framework to the European continent. It encompasses Spain and Portugal, offering localized insights into climate variability. Most recently, Garrido et al. (2023) adapt the ACI methodology to French data, creating the French Actuarial Climate Index. The continued adoption of the methodology worldwide demonstrates a growing global interest in effectively measuring and understanding climate risks.
@@
This study aims to compare the Actuaries Climate IndexTM, with long-established weather-based indexes frequently used in energy and derivatives markets, by using regression models: generalized linear model (GLM) and generalized additive model (GAM), as well as machine learning algorithms: extreme gradient boosting (XGB) and light gradient boosting machine (LGBM). By evaluating these indexes based on their performance in predicting crop yields, a critical factor in the agricultural sector, the study demonstrates the robustness of the ACI, showcasing its potential for real-time applications and broader applicability across various sectors. Existing literature demonstrates the applicability of the ACI in various fields, particularly insurance and finance. Jiang and Weng (2019) analyze ACI’s predictive power for stock returns of agriculture-related companies in the U.S. and Canada, concluding that ACI could support a profitable trading strategy. Pan et al. (2022) assess ACI’s effectiveness in predicting corn yields in the Midwest region of the U.S. using linear and probit regression models, finding satisfactory predictive power. Zhou and Vilar-Zanón (2024) explore the relationship between the Spanish Actuarial Climate Index (SACI) and hailstorm-related losses in the wine grape insurance industry through linear and quantile regression models, noting a potential rise in monthly losses corresponding to future SACI increases.
@@
Beyond comparing the predictive power of the indexes, this paper investigates the individual contribution of each weather index to model performance by fitting 22 distinct models, each constructed with different explanatory variables. This structured approach allows for a detailed assessment of how the inclusion of diverse weather indexes enhances the models’ predictive accuracy and its explanatory power. Since all explanatory variables are weather-based, they may share common weather driven patterns, potentially leading to multicollinearity issues in regression models. Additionally, achieving comparable model performance with fewer explanatory variables is both practical and desirable. To address these challenges, dimension reduction techniques specifically, principal component analysis (PCA) and functional principal component analysis (FPCA) are applied to identify the most influential variables in the data sets. This approach not only mitigates multicollinearity issues but also simplifies the models, improving efficiency and interpretability.
@@
The outcomes indicate that both ACI components and weather-based indexes exhibit robust predictive capabilities for crop yields (corn, wheat, soybeans) in the U.S. market. Furthermore, the findings reveal that incorporating indexes representing additional weather conditions—such as drought, wind, and sea-level—alongside traditional factors like air temperature and precipitation, substantially enhances model performance. This highlights the importance of a comprehensive approach that considers a broader spectrum of climatic variables to improve the accuracy and reliability of crop yield predictions.
@@
The paper is organized as follows: Section 2 summarizes the climate indexes used in the study. Section 3 describes the proposed two-stage methodology and presents the framework of the 22 models developed for crop yield prediction. The properties of the data and the results of the analyses are provided in Section 4. Finally, Section 5 concludes the paper with a summary of the key findings and their implications. Additionally, regression models, machine learning algorithms and dimensionality reduction techniques used in the analyses are reviewed in the Appendix A.
@@ 2 Climate Indexes
We focus on weather-based indexes (WBI) such as Cooling Degree Day (CDD), Heating Degree Day (HDD) and Cumulative Rainfall Index (PRE), which are directly tied to temperature and rainfall. These indexes are particularly advantageous due to their reliance on single weather variables, making them simple to understand, straightforward to compute and convenient for comparative analysis with the Actuaries Climate IndexTM. Detailed formulations of these weather-based indexes and outlines off the corresponding components within the ACI framework are given below.
@@ 2.1 Weather-Based Indexes
CDD and HDD are temperature-based indexes, with CDD capturing warm weather conditions and HDD representing cool weather conditions. PRE, on the other hand, quantifies rainfall amounts. These indexes are initially calculated for daily time periods, and their values for longer intervals, such as monthly or meteorological seasons, are derived by summing the daily data. Briefly, CDD measures how much the outside air temperature exceeds a base level temperature (usually taken as 65°F or 18°C), and HDD measures the opposite, the temperature falls below a base level (Jewson and Brix 2005).
Let $T_{i}^{\text{max}}$ and $T_{i}^{\text{min}}$ denote the maximum and minimum temperatures measured on day $i$. The average temperature for day $i$ is defined as,
@@
If the average temperature is higher than 65°F for CDD, and lower than 65°F for HDD, the difference becomes the $CDD_{i}$ and $HDD_{i}$ value for that day, respectively, given by,
$CDD_{i}=max\{T_{i}-65,0\},$ (2)
$HDD_{i}=max\{65-T_{i},0\}.$ (3)
On the other hand, PRE represents the total volume of rainfall accumulated over a specified time period (Odening et al. 2007), and is calculated by summing all daily rainfall measurements, expressed in millimeters, for the chosen time frame, such as a month or a meteorological season, described as
$PRE=\sum_{t=1}^{N}r_{t},$ (4)
where $r_{t}$ denotes the daily rainfall amount on day $t$, and $N$ is the total number of days in the specified period.
@@ 2.2 Actuaries Climate IndexTM
The ACI is designed to improve understanding of climate trends and their potential impact, incorporating various climate variables that signify extreme weather conditions. It comprises six components: high temperature (T90), low temperature (T10), precipitation (P), drought (D), wind speed (W), and sea-level (S). Since each component reflects different weather conditions, the ACI uses standardized values for all components to integrate them into a single index. Standardization is based on the mean and standard deviation of each component during the designated reference period of 1961 to 1990 (ACI 2018).
The components of the ACI used in our comparison analyses are briefly defined as follows:
@@
The standardized values of each component are computed both on a monthly and seasonal basis. As described in Zhou et al. (2023), for each month ($j$) and year ($k$), these are:
$T90_{STD}(j,k)=\frac{T90(j,k)-\mu_{T90(j)}}{\sigma_{T90(j)}},$ (5)
$T10_{STD}(j,k)=\frac{T10(j,k)-\mu_{T10(j)}}{\sigma_{T10(j)}},$ (6)
$P_{STD}(j,k)=\frac{Rx5day(j,k)-\mu_{Rx5day(j)}}{\sigma_{Rx5day(j)}}.$ (7)
Here, $\mu_{\cdot(\cdot)}$ and $\sigma_{\cdot(\cdot)}$ are the mean and standard deviation of corresponding components for the reference period, respectively. The composite ACI is derived by averaging the standardized values of all six components, with the T10 component being assigned a negative value to account for its inverse relationship with warm weather conditions, and expressed as:
@@
Detailed information of ACI can be found in ACI (2018).
3 Methodology
The primary objective of this study is to assess the robustness of the ACI as a climate impact measure in the agriculture sector. Given that ACI is a relatively recent addition to the literature, its performance in predicting crop yields is compared with the performance of long-established weather-based indexes. By evaluating the predictive capabilities of climate indexes, the study seeks to demonstrate ACI’s potential for real-time applications and its broader utility across various sectors.
The analyses are conducted in two stages. The first stage seeks similarities between WBI and ACI components. To do so, we first check how strongly correlated these two indexes are and define their empirical cumulative distribution functions. The assessment is conducted across three groups, each representing different weather conditions, enabling a systematic and detailed comparison of the two types of indexes, WBI and ACI. The second stage focuses on testing and comparing the predictive performance of climate indexes on crop yields. Crop yields are predicted by using regression models and machine learning algorithms that comprising a total of 22 distinct models, each constructed with different explanatory variables. This approach allows for a detailed analysis of the predictive performance of WBI and ACI, ensuring that variations in their predictive capabilities are systematically evaluated.
3.1 First Stage
Climate indexes that measure similar weather conditions are grouped into categories to facilitate meaningful comparisons. The grouping structure is based on the specific meteorological conditions represented by each index, such as warm weather, cool weather, and precipitation. Despite differences in their calculation methods, CDD defined in equation 2 and T90 in 5 measure warm weather conditions, HDD in 3 and T10 in 6 capture cool weather conditions, and PRE in 4 and P in 7 represent precipitation severity. This grouping ensures a systematic and logical comparison between the indexes, and its structure is summarized in Table 1.
Table 1: Structure of groups
Group
Weather-Based
Indexes (WBI)
Components
of the ACI
Meteorological
Season
Measurement
1 CDD T90 Summer Warm weather
2 HDD T10 Winter Cool weather
3 PRE P Spring Precipitation
For the comparisons, meteorological seasonal values (three months ending February, May, August, November) are used instead of monthly values. This approach aligns with the practical applications of climate indexes. For instance, CDD is commonly used to estimate energy demand during summer months, particularly for cooling, while HDD is used to evaluate heating needs during winter months (Spinoni et al. 2018). Regarding precipitation, its timing plays a critical role in agricultural productivity. Wang et al. (2016) highlight that wheat yields are especially sensitive to rainfall during the rainy season, while Peltonen-Sainio et al. (2011) point out that precipitation before harvest often negatively impacts crop yields. Therefore, in the precipitation group, index comparisons are performed for the spring months to encompass the rainy season preceding the general harvest period, which typically occurs during the summer months for most crops.
The dependence between the indexes are analyzed using the Spearman rank correlation coefficient (Zar 2005), a non-parametric measure that evaluates the strength and direction of the monotonic relationship between two variables. Distribution comparisons are performed using the Kolmogorov-Smirnov (K-S) test (Massey Jr 1951), which evaluates the maximum difference between the empirical cumulative distribution function values of the two data sets.
3.2 Second Stage

### v2

@@ Abstract
Climate change poses significant challenges to the agricultural and financial sectors, affecting crop productivity and overall financial stability. This study evaluates the robustness of the Actuaries Climate IndexTM (ACI), a newer entrant in the field as a tool for measuring climate impacts, by comparing its explanatory power with well-established weather-based indexes (WBIs) across two key sectors. In the agricultural context, the yields of three major crops are predicted using generalized statistical models and advanced machine learning algorithms with climate indexes as explanatory variables. To enhance model reliability and address multicollinearity among weather-related variables, the study also incorporates both principal component analysis and functional principal component analysis. A total of 22 models, each constructed with different sets of explanatory variables, demonstrate the significant impact of wind speed and sea-level changes, alongside temperature and precipitation, on crop yield variability across six regions of the United States. For the financial market application, the analysis adapts the weather-derivative framework, as it is a critical instrument for energy companies, insurers, and agribusinesses seeking to hedge against weather-related risks. By analyzing the payoffs of derivative contracts that use WBIs and ACI components as underlying variables, the findings reveal that the ACI framework holds a strong potential as a comprehensive climate risk indicator, not only for the agricultural sector but also for the finance and insurance industries.
Keywords actuaries climate index $\cdot$ weather-derivative pricing $\cdot$ crop yield prediction $\cdot$ weather-based indexes $\cdot$ machine learning
@@ 1 Introduction
Climate change refers to long-term shifts in temperature and weather. Recent reports from the Intergovernmental Panel on Climate Change highlight the significant increase in global temperatures over the past few decades (IPCC 2022; Masson-Delmotte et al. 2021). The consequences of this warming extend across multiple domains, adversely affecting human health, global food security, financial stability, and the insurance industry. From a public health perspective, WHO (2014) project that climate change could lead to an additional 250,000 deaths per year between 2030 and 2050. In the financial sector, Gall et al. (2011) report that direct losses from natural hazards in the United States (U.S.) are steadily increasing and may reach between \$300 and \$400 billion within a single decade, driven by the increasing frequency and severity of extreme weather events. In terms of global food security, empirical studies by Lobell and Gourdji (2012), Hu et al. (2024), and Adhikari et al. (2015) demonstrate that the yields of major crops are significantly affected negatively by increasing temperatures and changing precipitation patterns. Similarly, within the insurance industry, Zhou et al. (2024) discuss the profound impact of climate change across various insurance sectors, underscoring the urgency for standardized methodologies to measure and quantify climate risks.
Climate impact is mostly measured and interpreted by indexes. In the literature, various indexes such as the Climate Extremes Index (Karl et al. 1996), Greenhouse Climate Response Index (Karl et al. 1996), Palmer Drought Severity Index (Palmer 1965), Cooling Degree Days (Jewson and Brix 2005), Heating Degree Days (Jewson and Brix 2005) and Cumulative Rainfall Index (Odening et al. 2007) are developed to quantify climate change, enabling researchers and policymakers to assess systematically its impact. In recent years, the Actuaries Climate IndexTM (ACI) emerges as a pivotal tool for understanding and monitoring climate risks, offering valuable insights for insurance companies, governments, and the general public. Developed in 2016 by the four major North American actuarial societies, the ACI (2018) uses historical meteorological data from the U.S. and Canada to track extreme conditions in air temperature, precipitation, drought, wind, and sea-level rise.
@@
The practical methodology of the ACI is quickly adapted for use in other countries. Curry (2015) demonstrates that the ACI framework is effectively applicable to regions like the United Kingdom and Europe. The Australian Actuaries Climate Index (AACI 2018), which is tailored to Australian meteorological conditions, is introduced, while Nevruz et al. (2022) develop the Turkish Actuaries Climate Index, specifically for the Ankara region in Türkiye. Similarly, Zhou et al. (2023) launch the Iberian Actuarial Climate Index, encompassing Spain and Portugal, marking the first application of the ACI framework to the European continent. Most recently, Garrido et al. (2023) and Rogo et al. (2025) extend this framework to France and Italy, respectively, creating the French and Italian Actuarial Climate Indexes. The continued international adoption of the ACI methodology demonstrates a growing global interest to quantifying, monitoring, and managing climate risks through standardized and comparable metrics.
@@
Several studies in the literature also demonstrate the applicability of the ACI in various fields, particularly in insurance and finance. Jiang and Weng (2019) analyze ACI’s predictive power for stock returns of agriculture-related companies in the U.S. and Canada, concluding that ACI could support a profitable trading strategy. Pan et al. (2022) assess ACI’s effectiveness in predicting corn yields in the Midwest region of the U.S. using linear and probit regression models, finding satisfactory predictive power. Zhou and Vilar-Zanón (2024) explore the relationship between the Spanish Actuarial Climate Index and hailstorm-related insurance losses for grapes grown for wine producing, through linear and quantile regression models.
However, given its relatively recent addition, there remains a notable gap in the literature where a systematic evaluation of the ACI’s predictive capability against traditional weather indexes would provide a significant contribution, thereby enhancing its applicability across broader fields and sectors. By comparing the explanatory power of the ACI with well-established weather-based indexes (WBIs) commonly used in the energy and derivatives markets, this study aims to evaluate the robustness of the ACI as a tool for measuring climate impacts in both the agricultural sector and financial markets. In the agricultural context, weather-based indexes are used to predict the yields of three major crops using regression models, namely Generalized Linear Model (GLM) and Generalized Additive Model (GAM), alongside advanced machine learning algorithms such as Extreme Gradient Boosting (XGB) and Light Gradient Boosting Machine (LGBM). For the financial market application, the analysis focuses on the payoffs of weather derivative contracts that use WBIs and ACI components as underlying variables. To the best of our knowledge, beyond comparing the ACI with traditional weather indexes, this study makes an additional novel contribution by presenting the first application of the ACI within a weather-derivative framework which is critical instrument for energy companies, insurers, and agribusinesses seeking effective hedging mechanisms against weather-related risks.
@@
Regarding crop yield prediction, this study also investigates the individual contribution of each weather index to model performance by fitting 22 distinct models, each constructed with different sets of explanatory variables. Since all explanatory variables are weather-based, they may share common weather driven patterns, potentially leading to multicollinearity issues in regression models. Additionally, achieving comparable model performance with fewer explanatory variables is both practical and desirable. To address these challenges, dimension reduction techniques specifically, Principal Component Analysis (PCA) and Functional Principal Component Analysis (FPCA) are applied to identify the most influential variables in the data sets. This approach not only mitigates multicollinearity issues but also simplifies the models, improving efficiency and interpretability.
@@
The findings of this paper demonstrate that both ACI components and traditional weather-based indexes exhibit robust predictive capabilities for crop yields. Furthermore, the inclusion of additional weather conditions such as drought, wind, and sea-level alongside conventional variables like air temperature and precipitation, substantially enhances model performance. Additionally, the high degree of consistency observed among the derivative payoffs underscores the ACI framework’s potential as a comprehensive climate risk indicator, applicable not only to the agricultural sector but also to the finance and insurance industries.
@@
The paper is organized as follows: Section 2 summarizes the climate indexes, whereas Section 3 describes the proposed methodology for predicting crop yields and pricing weather derivatives based on these indexes. The properties of the data and the results of the analyses are provided in Section 4. Finally, Section 5 concludes the paper by summarizing the key findings and discussing their broader implications. Additional information on the regression models, machine learning algorithms, and dimensionality reduction techniques used in the analyses are provided in the Appendix A.
@@ 2 Climate Indexes
We focus on weather-based indexes, specifically Cooling Degree Days (CDD), Heating Degree Days (HDD), and the Cumulative Rainfall Index (PRE), which are directly linked to temperature and precipitation. These indexes hold particular advantages because they rely on a single weather variable, making them not only simple to interpret but also straightforward to compute. Their relative ease of use makes them especially suitable for comparative analyses with the components of the Actuaries Climate IndexTM. In the following subsections, the detailed formulations of these weather-based indexes and outlines of the corresponding components within the ACI framework are presented.
@@ 2.1 Weather-Based Indexes
CDD and HDD are temperature-based indexes, with CDD capturing warm weather conditions and HDD representing cool weather conditions. PRE, on the other hand, quantifies rainfall amounts. These indexes are initially calculated on a daily basis, and their values for longer intervals, such as monthly or meteorological seasons, are derived by summing the daily data. Briefly, CDD measures how much the daily average outside air temperature exceeds a base level temperature (usually taken as 65°F or 18°C), and HDD measures the opposite, the temperature falls below a base level (Jewson and Brix 2005).
Let $T_{i}^{\text{max}}$ and $T_{i}^{\text{min}}$ denote the maximum and minimum temperatures measured on day $i$. The average temperature for day $i$ is defined as:
@@
If the average temperature is higher than 65°F for CDD and lower than 65°F for HDD, the differences for that day are calculated by:
$CDD=max\{T_{i}-65,0\},$ (2)
$HDD=max\{65-T_{i},0\},$ (3)
respectively. Additionally, PRE represents the total volume of rainfall accumulated over a specified time period and is calculated by summing all daily rainfall measurements, expressed in millimeters (Odening et al. 2007):
$PRE=\sum_{i=1}^{L}r_{i},$ (4)
where $r_{i}$ denotes the daily total rainfall amount on day $i$, and $L$ is the total number of days in the chosen time frame, such as a month or a meteorological season.
@@ 2.2 Actuaries Climate IndexTM
The ACI is designed to improve understanding of climate trends and their potential impact, incorporating various climate variables that signify extreme weather conditions. It comprises six components: high temperature (T90), low temperature (T10), precipitation (P), drought (D), wind speed (W), and sea-level (S). Since each component reflects different weather conditions, the ACI uses standardized values (anomalies) for all components to integrate them into a single index. Standardization is based on the mean and standard deviation of each component during the designated reference period of 1961 to 1990 (ACI 2018). The components of the ACI used in the analyses are defined as follows:
@@
The standardized values of each component (anomalies) are computed both on a monthly and seasonal basis. For each month ($j$) and year ($q$), they are expressed as (Zhou et al. 2023):
$T90_{STD}(j,q)=\frac{T90(j,q)-\mu_{T90(j)}}{\sigma_{T90(j)}},$ (5)
$T10_{STD}(j,q)=\frac{T10(j,q)-\mu_{T10(j)}}{\sigma_{T10(j)}},$ (6)
$P_{STD}(j,q)=\frac{Rx5day(j,q)-\mu_{Rx5day(j)}}{\sigma_{Rx5day(j)}}.$ (7)
Here, $\mu_{\cdot(\cdot)}$ and $\sigma_{\cdot(\cdot)}$ are the mean and standard deviation of corresponding components for the reference period, respectively. The composite ACI is derived by averaging the standardized values of all six components, with the T10 component being assigned a negative value to account for its inverse relationship with warm weather conditions, and expressed as (ACI 2018):
@@
3 Proposed Methodology
Based on the main objectives, to evaluate the robustness of the ACI as a tool for measuring climate impacts in both the agricultural sector and financial markets, the study examines the performance of the ACI in two distinct applications: crop yield prediction as an agricultural case and the pricing of weather derivatives as a financial market case. Since the ACI is a relatively recent in the literature, its performance is compared against that of long-established weather-based indexes. By assessing the explanatory power of climate indexes, this study is expected to underscore the potential of the ACI for real-time applications and demonstrate its broader utility across diverse sectors.
3.1 Crop Yield Prediction